# Position: AI agents should be regulated based on autonomous action sequences

### v1

@@ 3 Alternative views
One broadly applicable metric for regulating AI agents is the amount of computation they require. Here, we explore how existing regulations and recommendations from AI researchers often center on computational resources. While these efforts primarily address the computational resources for training, we expand the discussion to include computational demands for inference. We then argue that limiting the focus to these aspects alone is inadequate for the existential risks associated with advanced AI agents.
3.1 Computational resources for development
In this section, we briefly review existing regulations, focusing on the EU AI Act and President Biden’s executive order, along with key recommendations from AI researchers.
3.1.1 EU AI Act
@@
European Parliament and Council of the European Union (2024) has established a set of rules (the EU AI Act) for the development, deployment, and use of AI within European Union. Its Chapter 5 defines “general-purpose AI models with systemic risk” and lists obligations for providers of such models. Here, general-purpose AI models essentially refer to FMs, which are pre-trained with self-supervised learning and can (be adapted to) perform a wide range of downstream tasks (see Article 3(63)). Also, systemic risk refers to “a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact … on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain” (see Article 3(65)).
In particular, a “general-purpose AI model shall be presumed to have high impact capabilities … when the cumulative amount of computation used for its training measured in [FLOPs] is greater than $10^{25}$” (see Article 51(2)). When this or other specified conditions are met, the provider of a general-purpose AI model is required to fulfill certain obligations, such as providing technical documentation about the model.
The advanced AI agent that we consider can certainly be classified as general-purpose AI. Also, the existential risks that we consider can be considered as systemic risks, although human extinction and irreversible global catastrophes are not discussed in the EU AI Act.
3.1.2 Biden’s executive order and related bills
In October 2023, Joe Biden, then president of the US, signed the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence Biden (2023) [^3] This executive order was repealed by President Trump.. Its Section 4.2 is dedicated to ensuring safe and reliable AI. In particular, it requires companies to report on “any model that was trained using a quantity of computing power greater than $10^{26}$ integer or [FLOPs]” until a set of technical conditions for models are defined by specified authorities (see Section 4.2(b)).
In this executive order, particular attention is paid to a dual-use FM, which refers to a FM that exhibits “high levels of performance at tasks that pose a serious risk to security, national economic security, national public health or safety, or any combination of those matters, such as by … permitting the evasion of human control or oversight through means of deception or obfuscation” (see Section 3(k)).
Following this executive order, almost 700 AI-related bills are introduced in 45 states across the United States in 2024 The Business Software Alliance (2024). A particularly interesting one is California Senate Bill 1047 (Safe and Secure Innovation for Frontier Artificial Intelligence Models Act) Wiener et al. (2024) [^4] The bill had passed the state legislature but was later vetoed by the Governor. The technical feasibility of the requirements has also been questioned by the community AI Alliance (2024).. Its Chapter 22.6 is devoted to safe and secure innovation for frontier artificial intelligence models, which cover “[a]n artificial intelligence model trained using a quantity of computing power greater than $10^{26}$ integer or floating-point operations.” In particular, the senate bill requires that “[b]efore beginning to initially train a covered model, the developer shall … [i]mplement the capability to promptly enact a full shutdown,” which completely halts the operations of the model.
The necessity of such an off-switch Hadfield-Menell et al. (2017) is motivated to prevent “critical harms,” which include “[m]ass casualties or at least five hundred million dollars (\$500,000,000) of damage resulting from an artificial intelligence model engaging in conduct that … [a]cts with limited human oversight, intervention, or supervision.”
3.1.3 Recommendations by scientists
We have seen that the computational resources used for training is one of the major criteria to judge whether an AI model can have unacceptably high risk. Although specific values of the threshold such as $10^{25}$ or $10^{26}$ FLOPs are subject to change, the amount of computation for training is one of the most reliable metrics that AI researchers can currently provide for approximating the performance and potential risks of a wide range of FMs.
@@ 3.1.3 Recommendations by scientists
For example, Anderljung et al. (2023) recommend to identify sufficiently dangerous frontier AI models on the basis of whether they are trained with more than $10^{26}$ FLOPs of computation. More recently, Cohen et al. (2024) argue that “[s]ystems should be considered ‘dangerously capable’ if they are trained with enough resources to potentially exhibit those dangerous capabilities, and regulators should not permit the development of dangerously capable LTPAs.” Although they do not specify exactly what is considered as enough resources, the amount of computation for training is the only specific criterion that they suggest to determine whether an LTPA can have existential risk.
@@
Future of Life Institute (2023b) have also provided recommendations for governments on managing AI risks. The recommendations include mechanisms such as auditing, certification, and regulation, grounded in the assumption that “[t]he amount of compute used to train a general-purpose system largely correlates with … the magnitude of its risks” Future of Life Institute (2023b). These recommendations have been made by following an open letter that called for “all AI labs to immediately pause for at least 6 months the training of AI systems more powerful than GPT-4,” which was issued in response to the severe societal risks posed by advanced AI systems and signed by AI researchers Future of Life Institute (2023a).
@@
In fact, the amount of computation used for training may serve as a sufficiently reliable predictor of the performance and risks of today’s LLMs. This is because most existing LLMs share the same Transformer architecture, differing primarily in their size and the volume of training data. Several studies have examined scaling laws that describe the relationship between a model’s optimal size, the amount of training data, and the computational budget required Kaplan et al. (2020); Hoffmann et al. (2022). Research has also shown that various abilities, such as multi-step reasoning, tend to emerge as the computational resources used for training increase Wei et al. (2022a).
@@
These scaling laws can, in turn, be used to estimate the FLOPs needed to train existing LLMs. For example, Anil et al. (2023) propose a heuristic suggesting that an LLM should be trained with $6\,N\,D$ FLOPs, where $D$ is the amount of training data, and $N$ is the model size. Using this heuristics, Llama 3.1 405B trained on 15 trillion tokens is estimated to require approximately $4\times 10^{25}$ FLOPs—an amount closely aligning with the thresholds specified in the EU AI Act and President Biden’s executive order.
3.2 Computational resources for operation
We have seen that the amount of computation used for training is one of the key criteria that is used today to regulate highly capable FMs. Although such regulations may be effective for AI agents whose capabilities primarily and directly stem from traditional FMs, they obviously fail to regulate advanced AI agents that gain substantial reasoning capabilities from computation at inference time.
@@ 3.2 Computational resources for operation
A natural approach is to regulate advanced AI agents based on the amount of computation used not only for training but also for inference. While this approach may be effective for some AI agents, we argue that it is insufficient for advanced AI agents, at least for the following four reasons.
@@
First, scaling laws for inference-time computation are much less established than those for training-time computation. Recently, scaling laws have been studied for a limited number of inference strategies in LLMs Chen et al. (2024); Brown et al. (2024); Snell et al. (2024); Wu et al. (2024). However, there are many potential inference strategies, and their performance can scale very differently. The scaling law for inference also depends on how difficult the task is and what LLM is used Snell et al. (2024); OpenAI (2024).
@@
Second, inference can be performed in parallel by multiple entities. For example, several entities may operate the same AI agent that perform reasoning with MCTS Luo et al. (2024); Zhang et al. (2024b), either collaboratively or independently, possibly without knowing each other. Since MCTS is a randomized algorithm, the likelihood that one of the agents optimally solves the task increases with the number of agents. However, this also means that this agent may exploit a loophole, solving the task super-optimally in a way that violates critical constraints, potentially leading to catastrophic outcomes.
@@
Third, it is not always clear what constitutes a single run of inference. For instance, the results of one inference run may be stored and used in another. Intermediate results could also be stored and later retrieved by a different AI agent, who may or may not be aware that the information is from a previous inference run. More broadly, reasoning can be enhanced through retrieval augmentation Pouplin et al. (2024), where retrieved information may have been generated with substantial computational resources.
@@
Finally, AI agents may continually learn over time, which makes it difficult to separate inference from training. For example, an AI agent may perform reasoning with MCTS, with an LLM performing a step in the process. Once the agent identifies a good sequence of steps, it may fine-tune the LLM in a way that the LLM can perform the entire sequence in a single step, bypassing the reasoning process. As this learning progresses, the agent will gain the ability to perform high-level reasoning with limited computation.
4 Approach based on action sequence
We have seen that the amount of computation is one of the best criteria that we can use today to regulate advanced AI agents. Although we argue that the amount of computation alone is insufficient, we do not mean to dismiss it. We should regulate advanced AI agents from various perspectives to mitigate existential risks. Currently, we lack technologies that allow us to regulate advanced AI agents more effectively than or complementary to the amount of computation.
@@ 4 Approach based on action sequence
Here, we suggest a potential approach as a basis of discussion. The suggested approach by no means solves the problem of existential risk. Rather, it serves as a baseline and a call for action for further research on this matter. We hope that the community will build upon it or develop entirely different approaches, eventually leading to technologies that can properly govern and regulate advanced AI agents.
@@
In the following, we say that the amount of computation, an AI agent, or its action sequence is acceptable when it involves low existential risk. Also, we say that any of the three is strongly acceptable when it involves very low existential risk, which we will slightly more formalize in the sequel and in Appendix B.1.
@@ 4.1 Why and why not the amount of computation?
We first discuss, in a slightly formal manner, why and when the amount of computation may or may not work as a criteria of whether an AI agent is acceptable. Consider the process of building an AI agent $G$ and letting $G$ generate and take actions (Figure 1). The $G$ is trained with a given configuration $C$, which includes training data, the architecture of a neural network, the amount of computation to use, and other conditions of a training algorithm. Such a trained $G=f(C)$ generates an action $A$ (e.g., a document or a command to use an external tool), possibly after searching optimal actions with MCTS. The $A$ is then taken in a cyber-physical world, and $G$ perceives a part the world $x$ (e.g., getting a prompt) to generate further actions depending on $x$. The $G$ typically determines the conditional distribution of $A$ given $x$. Namely, $A$ is a sample from the conditional distribution, which is determined by $G$ that is trained with $C$: $A\sim g(G,x)$, where $G=f(C)$.
When we say that an AI agent trained with the amount of computation $C_{1}$ is acceptable, it is assumed that the acceptability does not decrease much when we only slightly increase the amount of computation. More formally, suppose that we empirically know that any AI agent is strongly acceptable as long as it is trained with at most $C_{1}^{\prime}$ FLOPs of computation (e.g. $C_{1}^{\prime}=10^{24}$). [^5] For example, an AI agent may be considered strongly acceptable if it does not show any symptoms of unacceptable behaviors for a certain period of time. Based on the behaviors of existing LLMs, we may say that an LLM trained with at most $10^{24}$ FLOPs of computation is strongly acceptable. Then the assumption is that the AI agent trained with $C=(C_{1},C_{-1})$ is acceptable regardless of $C_{-1}$ as long as $d(C_{1},C_{1}^{\prime})\leq\varepsilon$ for some metric $d$ and a threshold $\varepsilon$ (e.g. $C_{1}\leq 10^{25}$ if $\varepsilon=1$ and $d$ is the difference in $\log_{10}$ of FLOPs used for training).
While this assumption may be a good approximation for AI agents that are primarily based on LLMs with standard Transformer architectures whose performance is well characterized by scaling laws, advanced AI agents with reasoning and planning capabilities are unlikely to satisfy the assumption, since they would heavily rely on inference-time computation. Also, the assumption may no longer hold even for LLMs when new technologies are introduced to enable training an LLM having similar capabilities with substantially less computational resources.
@@ 4.2 Similarity between action sequences
As is evident from Figure 1, it is the action that makes direct impacts on the environment and can pose existential risks. Hence, rather than regulating configurations, we may directly regulate the actions taken by AI agents. In Appendix A, we will also discuss an alternative approach of regulating AI agents, particularly based on their output distributions, rather than their actions or configurations.
However, whether an action is acceptable or not typically depends on when, where, and whom the action is taken (i.e., the state of the environment). In fact, the action space of an AI agent can typically be designed by humans, who could ensure that each individual action is acceptable at normal states. For example, each action may generate a document, search information from the Web, use an external tool via API, or other elementary steps that would look acceptable when they are considered individually. It is the sophisticated sequence of actions that can gradually transform the state into unstable one, which can then make the next action unacceptable.
Hence, in the study of the safety of AI agents and RL agents, it is often modeled that those agents operate in Markov decision processes (MDPs). Then some states are considered unsafe or some state-action pairs are considered unsafe. One does not say that some actions are unsafe, because those actions can be easily excluded from the action space.
@@
While these conventional approaches of AI/RL safety are principled in theory, they cannot be easily applied to the scenarios where it is hard to define, represent, or observe the (Markovian) state. AI agents may operate in open-ended real-world environments, and they observe only very partial information about the state of the environment. The concept of existential risks is rather vague, and we do not know specifically what can cause irreversible global catastrophes or human extinction. This makes it infeasible to choose specific features of the environmental state to observe, but the entire state of the environment cannot be observed either.
Consider the states and actions in an MDP as a graphical model in Figure 2. We want to avoid taking an action $a_{T-1}$ that can lead to an unacceptable state $s_{T}$. Since the distribution of $s_{T}$ is determined by the pair $(s_{T-1},a_{T-1})$ according to a transition probability $p(s_{T}|s_{T-1},a_{T-1})$, it makes sense to consider the safety of the pair $(s_{T-1},a_{T-1})$ as in the conventional approaches.
Figure 2: States and actions in a Markov decision process
@@
The distribution of $s_{T}$ can also be recursively computed (i.e. $p(s_{t}|s_{t-1},a_{t-1})$ for $t=1,\ldots,T$) from the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$. Although the lack of knowledge about $s_{T-1}$ makes $s_{T}$ less predictable, we may still consider whether the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, or whether $a_{T-1}$ is acceptable given $(s_{0},a_{0},\ldots,a_{T-2})$. This, however, still requires knowing the initial state $s_{0}$, which cannot be observed in general.
@@
Our proposal is to study whether the action-sequence $(a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable. This may correspond to ensuring that the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable for any $s_{0}$. Although this may sound overly conservative, the dependence on $s_{0}$ may not be significant in practice. For example, an ergodic MDP gets close to the steady state distribution after a mixing time, regardless of the initial state. Although ergodicity generally does not hold in real-world environments, we expect that the action-sequence is a practical compromise to define a measure of acceptability.
Similar to the amount of computation, we may say that an action-sequence is acceptable when it is sufficiently close to the action-sequence that is known to be strongly acceptable. For example, if $(a_{0},\ldots,a_{T-1})$ is strongly acceptable, $(a_{0},\ldots,a_{T-1},a_{T})$ would be acceptable for any $a_{T}$. Note that an action is different from a reasoning step but rather an operation that has some impacts on the environment. For example, an action may be an API (Application Programming Interface) call that is selected and executed after a heavy reasoning process. In Appendix B, we discuss how the strongly acceptable set may be gradually expanded, in relation to safe exploration in RL.
@@
There are many ways to measure the distance between sequences. Dynamic time warping is a representative method. However, a simple measure such as the difference in the length of the action sequence may be more widely applicable. Specifically, we may say that the action-sequence of length $T$ is acceptable after getting sufficient empirical evidence that any action-sequences of length $T-1$ is strongly acceptable, regardless of the initial state. Then AI agents are allowed to autonomously take actions for at most $T$ consecutive steps. This simplicity is reminiscent of the amount of computation, where we say that the FM trained with $10^{25}$ FLOPs is acceptable, since we have sufficient evidence that FMs trained with $10^{24}$ FLOPs are strongly acceptable, regardless of how they are developed or operated.
Note that our proposal is to regulate the sequence of actions taken autonomously by AI agents. This means that a new action sequence may start after a human intervention. Also, when an agent cannot accomplish its goal with $T$ actions, it may consult a human and take further actions under a careful supervision of the human.

### v2

@@ 3 Alternative views
One broadly applicable metric for regulating AI agents is the amount of computation they require. Here, we explore how existing regulations and recommendations from AI researchers often center on computational resources. While these efforts primarily address training compute, we expand the discussion to include inference compute. We then argue that limiting the focus to these aspects alone is inadequate for the existential risks associated with advanced AI agents.
3.1 Training computate
@@
The European Union (EU) has established a set of rules (EU AI Act) for the development, deployment, and use of AI within EU [40]. Its Chapter 5 defines “general-purpose AI models with systemic risk” and lists obligations for providers of such models. Here, general-purpose AI models essentially refer to FMs, which are pre-trained with self-supervised learning and can (be adapted to) perform a wide range of downstream tasks (see Article 3(63)). Also, systemic risk refers to “a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact … on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain” (see Article 3(65)). In particular, a “general-purpose AI model shall be presumed to have high impact capabilities … when the cumulative amount of computation used for its training measured in [FLOPs] is greater than $10^{25}$” (see Article 51(2)). When this or other specified conditions are met, the provider of a general-purpose AI model is required to fulfill certain obligations, such as providing technical documentation about the model.
While existential risks are not explicitly considered systemic risks under this definition, human extinction could still be relevant if it results from large-scale failures that affect public health, safety, security, fundamental rights, or society in a way that propagates across the AI value chain. Also, certain advanced AI agents could meet the systemic risk criteria in the EU AI Act due to their high-impact capabilities and widespread deployment. Specifically, an AI agent with advanced decision-making autonomy, broad adaptability, and deep integration into critical infrastructures could introduce large-scale disruptions or cascading failures, thus meeting the criteria of systemic risk.
Regulations centered on training compute were also a central focus of the executive order signed by then-President Biden in October 2023, as well as the AI-related legislative proposals that followed. Although this executive order was later repealed by President Trump, its influence on subsequent policy discussions remained significant. See Appendix A for details.
@@
Training compute is one of the most reliable metrics that AI researchers can currently provide for approximating the performance and potential risks of FMs. Anderljung et al.  recommend to identify sufficiently dangerous frontier AI models on the basis of whether they are trained with more than $10^{26}$ FLOPs of computation [6]. Cohen et al. argue that “[s]ystems should be considered ‘dangerously capable’ if they are trained with enough resources to potentially exhibit those dangerous capabilities, and regulators should not permit the development of dangerously capable LTPAs” [33]. Although they do not specify what constitute sufficient resources, training compute is the only specific criterion that they suggest to determine whether an LTPA can have existential risk.
@@
Future of Life Institute has also provided recommendations for governments on managing AI risks [45]. The recommendations include mechanisms such as auditing, certification, and regulation, grounded in the assumption that “[t]he amount of compute used to train a general-purpose system largely correlates with … the magnitude of its risks” [45]. These recommendations have been made by following an open letter that called for “all AI labs to immediately pause for at least 6 months the training of AI systems more powerful than GPT-4,” which was issued in response to the severe societal risks posed by advanced AI systems and signed by AI researchers and business leaders [44].
@@
Indeed, training compute serves as a sufficiently reliable predictor of the performance and risks of most existing LLMs. This is because they share the same Transformer architecture, differing primarily in their size and the volume of training data. Several studies have examined scaling laws that describe the relationship between a model’s optimal size, the amount of training data, and training compute [70, 58]. Research has also shown that various abilities, such as multi-step reasoning, tend to emerge as training compute increases [134].
@@
These scaling laws can, in turn, be used to estimate the FLOPs needed to train existing LLMs. For example, Anil et al.  propose a heuristic suggesting that an LLM should be trained with $6ND$ FLOPs, where $D$ is the amount of training data, and $N$ is the model size [7]. Using this heuristics, Llama 3.1 405B trained on 15 trillion tokens is estimated to require approximately $4\times 10^{25}$ FLOPs—an amount closely aligning with the thresholds recommended by researchers and specified in the EU AI Act.
Limitations of training compute have also been discussed in the literature [59, 74, 107]. In particular, thresholds such as $10^{25}$ and $10^{26}$ FLOPs are arbitrary and lack scientific grounding. Hooker argues that such thresholds should evolve with advances in data quality, optimization techniques, and architectures [59]. She also notes that training compute alone is insufficient, advocating for benchmarking AI agents against specific risks. However, when it comes to existential risks, we currently lack better evaluation methods than training compute.
3.2 Inference compute
@@
Although regulations based on training compute may be effective for agents whose capabilities primarily and directly stem from traditional FMs, they obviously fail to regulate those agents that gain substantial reasoning capabilities from inference compute. A natural approach is to regulate agents based also on inference compute. While this approach may be effective for some AI agents, we argue that it is insufficient for advanced AI agents, at least for the following four reasons.
@@
First, scaling laws for inference compute are far less established than those for training compute. Recent studies have explored inference scaling for a limited number of inference strategies in LLMs [31, 25, 116, 138], but many strategies remain unexamined and may scale differently. Moreover, inference scaling depends heavily on task difficulty and the specific LLM used [116, 94].
@@
Second, inference can be performed in parallel by multiple entities. For example, several entities may operate the same AI agent that performs reasoning with MCTS [85, 146], either collaboratively or independently, possibly without knowing each other. Since MCTS is a randomized algorithm, the likelihood that one of the agents optimally solves the task increases with the number of agents. However, it also increases the risk that one of the agents will exploit a loophole, solving the task super-optimally in a way that violates critical constraints, potentially leading to catastrophic outcomes.
@@
Third, it is not always clear what constitutes a single run of inference. For instance, the results of one inference run may be stored and used in another. Intermediate results could also be stored and later retrieved by a different AI agent, who may or may not be aware that the information is from a previous inference run. More broadly, reasoning can be enhanced through retrieval augmentation [101], where retrieved information may have been generated with substantial computation.
@@
Finally, agents that learn continuously blur the line between inference and training. For example, an AI agent may perform reasoning with MCTS, with an LLM performing a step in the process. Once the agent identifies a good sequence of steps, it may fine-tune the LLM in a way that the LLM can perform the entire sequence in a single step, bypassing the reasoning process. As this learning progresses, the agent will gain the ability to perform high-level reasoning with limited computation.
4 Approach based on the extent of autonomous operations
@@
In this section, we suggest a potential approach as a basis for discussion on regulating advanced AI agents. Although we argue that training compute alone is insufficient, we do not mean to dismiss its value. Rather, it should be seen as one component within a broader, multi-faceted regulatory framework. The approach we propose does not fully solve the problem of existential risk. Instead, it serves as a baseline and a call to action for further research. We hope the community will build upon this foundation or develop entirely new approaches, ultimately leading to technologies capable of properly governing and regulating advanced AI agents.
@@
In the following, we say that the amount of computation, an AI agent, or its action sequence is acceptable when it involves low existential risk. Also, we say that any of the three is strongly acceptable when it involves very low existential risk, which we will slightly more formalize in the sequel and in Appendix C.
@@ 4.1 Why and why not the amount of computation?
We first discuss, in a slightly formal manner, why and when the amount of computation may or may not work as a criteria of whether an AI agent is acceptable. Consider the process of building an AI agent $G$ and letting $G$ generate and take actions (Figure 1). The $G$ is trained with a given configuration $C$, which includes training compute, training data, the architecture of a neural network, and other conditions of a training algorithm. Such a trained $G=f(C)$ generates an action $A$ (e.g., a document or a command to use an external tool), possibly after searching optimal actions with MCTS. The $A$ is then taken in a cyber-physical world, and $G$ perceives a part the world $x$ (e.g., getting a prompt) to generate further actions depending on $x$. The $G$ typically determines the conditional distribution of $A$ given $x$. Namely, $A$ is a sample from the conditional distribution, which is determined by $G$ that is trained with $C$: $A\sim g(G,x)$, where $G=f(C)$.
When we say that an agent built with training compute $C_{1}$ is acceptable, it is assumed that the acceptability does not decrease much when we only slightly increase training compute. More formally, suppose that we empirically know that any agent is strongly acceptable as long as it is built with at most $C_{1}^{\prime}$ training compute (e.g. $C_{1}^{\prime}=10^{24}$ FLOPs). [^3] For example, an agent may be considered strongly acceptable if it does not show any symptoms of unacceptable behaviors for a certain period of time. Based on the behaviors of existing LLMs, we may say that an LLM trained with at most $10^{24}$ FLOPs of computation is strongly acceptable. Then the assumption is that the agent trained with $C=(C_{1},C_{-1})$ is acceptable regardless of $C_{-1}$, configurations other than training compute, as long as $d(C_{1},C_{1}^{\prime})\leq\varepsilon$ for some metric $d$ and a threshold $\varepsilon$ (e.g. $C_{1}\leq 10^{25}$ FLOPs if $\varepsilon=1$ and $d$ is the difference in $\log_{10}$ of training compute in FLOPs).
While this assumption may reasonably hold for the agents based on LLMs with standard Transformer architectures—whose performance aligns well with scaling laws—it is less applicable to advanced AI agents with reasoning and planning capabilities, which depend heavily on inference compute. Moreover, the assumption breaks down even for LLMs when new technologies enable similar capabilities with significantly reduced training compute.
@@ 4.2 Similarity between action sequences
As is evident from Figure 1, it is the action that makes direct impacts on the environment and can pose existential risks. Hence, rather than regulating configurations, we should directly regulate the actions taken by agents. In Appendix B, we also discuss an alternative approach of regulating agents based on their output distributions, rather than their actions or configurations.
Since the action space of an agent is typically designed by humans, inherently dangerous actions—such as launching a nuclear weapon—can be explicitly excluded [^4] The U.S. and China have agreed that AI systems should not be granted control over nuclear weapons [106].. The action space should include only those actions that are considered acceptable when they are taken individually, at normal states. Examples include generating a document, searching information from the Web, using an external tool via API (Application Programming Interface), and other elementary steps.
However, the acceptability of an action often depends heavily on context—specifically, when, where, and to whom it is applied (i.e., the current state of the environment). A sophisticated sequence of otherwise acceptable actions can gradually shift the environment into an unstable or unsafe state, ultimately rendering subsequent actions—normally considered safe—unacceptable. To capture this dynamic, the safety of AI/RL agents is often studied within the framework of Markov Decision Processes (MDPs), where certain states or state-action pairs are explicitly designated as unsafe.
@@
While these conventional approaches of AI/RL safety are principled in theory, they are difficult to apply in scenarios where the (Markovian) state is hard to define, represent, or observe. Agents operate in open-ended, real-world environments where they observe only very partial information about the state. The concept of existential risks is rather vague—we lack precise knowledge of what might trigger irreversible global catastrophes or human extinction. This uncertainty makes it impractical to identify specific state features to monitor, yet observing the full state is equally infeasible.
@@
We should avoid taking the action $a_{T-1}$ that can lead to an unacceptable state $s_{T}$. Since the distribution of $s_{T}$ is determined by $(s_{T-1},a_{T-1})$ according to the transition probability $p(s_{T}|s_{T-1},a_{T-1})$, it makes sense to consider the safety of the pair $(s_{T-1},a_{T-1})$ as in the conventional approaches. The distribution of $s_{T}$ can also be recursively computed from the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ by applying $p(s_{t}|s_{t-1},a_{t-1})$ for $t=1,\ldots,T$. We may thus consider whether $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, or whether $a_{T-1}$ is acceptable given $(s_{0},a_{0},\ldots,a_{T-2})$. This, however, still requires knowing the initial state $s_{0}$, which cannot be fully observed.
@@
In this setting, our proposal is to study whether the action-sequence $(a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, independent of the initial state $s_{0}$. This corresponds to ensuring that the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable for any $s_{0}$ that we may normally encounter in the absent of the agent. Although this may appear overly conservative, the dependence on $s_{0}$ may not be significant in practice. For example, an ergodic MDP gets close to the steady state distribution after a mixing time, regardless of the initial state. Although ergodicity rarely hold in real-world environments, we argue that focusing on the action sequence offers a practical and tractable compromise for defining acceptability.
@@
To assess whether an action sequence is acceptable, we can compare it to a reference sequence known to be strongly acceptable. Various metrics exist for measuring sequence similarity. Dynamic time warping, for example, accounts for temporal misalignments, but simpler metrics—such as differences in sequence length—offer broader applicability, particularly in regulatory contexts. Specifically, we may say that the action-sequence of length $T$ is acceptable once there is sufficient empirical evidence that all sequences of length $T-1$ are strongly acceptable, regardless of the initial state. Then agents are allowed to autonomously take actions for up to $T$ consecutive steps. This simplicity is reminiscent of the approach with computational limit, where we say that the FM trained with $10^{25}$ FLOPs is acceptable, since we have sufficient evidence that FMs trained with $10^{24}$ FLOPs are strongly acceptable, regardless of how they are developed or operated.
There is flexibility in how the length of an action sequence is defined, which also depends on how the action space is defined. A straightforward approach is to count the number of actions in the sequence. However, this may not always be the most meaningful metric. For instance, one could instead focus on actions that have tangible impacts on the environment—such as API calls or interactions with physical systems including humans. In contrast, actions with negligible external effects—such as generating individual tokens or updating internal states during reasoning—may be too fine-grained to warrant consideration in the context of regulating existential risks. More generally, we may define a designated subspace of actions that are considered when evaluating the length of action sequences.
While action-sequence length is not a direct proxy for risk, allowing longer sequences generally expands the set of achievable outcomes for a given action space. Since usefulness and risk often scale together, an agent that can take longer sequences of actions is likely to exhibit more powerful—and potentially more hazardous—behaviors. In this sense, action length can be correlated with risk, particularly in the absence of additional safeguards. This does not imply that every longer sequence is more dangerous—but that the space of dangerous sequences grows with length.
Importantly, our proposal focuses on regulating the sequences of actions taken autonomously by AI agents. For instance, if an agent cannot achieve its objective within a limit of $T$ actions, it may consult a human and take further actions under a careful supervision of the human. In Appendix CD, we further discuss how the strongly acceptable set may be gradually expanded in connection with safe exploration in RL, and provide additional rationale for the similarity-based measures.